selbly decided today that he will save R15000 per quarter over the next 4 years.He will make the first deposit into a saving account in three months time and he will make his last deposit at the end of three years from now.how much will Selby have at the end of four years if interest is earned at 8.8% per annum ,compounded quarterly?
1
Expert's answer
2020-07-23T17:57:05-0400

Compounding quarterly means 4 times a year.

A = P (1 +r/100)n

1st deposit A = 15000 (1+r/100)n

quarterly in terms of years 4* 4 = 16

rate will be 8.8/4 =2.2

A = 15000 (1+ 2.2/100)16

A = 15000 ( 1.022)16

15000 * 1.4165

amount after 4 years =R 21247.50